A player hits a ball of mass 0.3kg which was moving eastwards with a velocity of 10m/s causing it now to move with a velocity of 15m/s westwards. The forces of the blow acts on for 0.01s . Calculate the average force exerted on the ball by the player.


1
Expert's answer
2021-11-07T19:28:00-0500

Let's choose eastwards as a positive direction. Then the change of the velocity was the following:


"\Delta v = 15m/s -(-10m/s) = 25m/s"

Accroding to the second Newton's law, the average force is:


"F_{av} = m\dfrac{\Delta v}{\Delta t}"

where "\Delta t = 0.01s" is the time the force acted on the ball, and "m=0.3kg" is its mass. Thus, obtain:


"F_{av} = 0.3kg\cdot \dfrac{25m/s}{0.01s} = 750 N"

Answer. 750 N.